OpenStudy (anonymous):

For the triangle shown, tan B = 3. (The triangle is not drawn to scale.) What is the value of x? Round your answer to the nearest tenth. A. 2.7 B. 5.0 C. 20.4 D. 24.0

OpenStudy (anonymous):



OpenStudy (anonymous):

@xVxNICKxVx @Bee_see

OpenStudy (ribhu):

Tan B = 8/x 3 = 8/x x = 8/3 = 2.66666666

OpenStudy (ribhu):

now u can round ur answer

OpenStudy (anonymous):

Thanks!